# Laser interferometry as a robust neuromorphic platform for machine learning

### v3

@@ Abstract
We present a method for implementing an optical neural network using only linear optical resources, namely field displacement and interferometry applied to coherent states of light. The nonlinearity required for learning in a neural network is realized via an encoding of the input into phase shifts allowing for far more straightforward experimental implementation compared to previous proposals for, and demonstrations of, in situ inference. Beyond in situ inference, the method enables in situ training by utilizing established techniques like parameter shift rules or physical backpropagation to extract gradients directly from measurements of the linear optical circuit. We also investigate the effect of photon losses and find the model to be very resilient to these.
journal: opticajournal
1 Introduction
@@
Modern deep learning models have reached unprecedented sizes with recent large language models having billions of trainable parameters and taking months of time for pretraining on clusters of graphics processing units (GPUs) [1, 2, 3]. The training and inference phases for these models require increasingly large amounts of energy due to the energy-efficiency limitations imposed by the modern silicon-based digital computers on which they are implemented, a phenomenon referred to as the ‘von Neumann bottleneck’. This bottleneck is a consequence of the separation of computing and memory into distinct units in the current von-Neuman computer architecture which necessitates costly data transfer between these units during computation [4, 5]. This challenge has motivated the development of neuromorphic computing, seeking hardware architectures inspired by the remarkable energy efficiency of the human brain [4, 5]. Consequently, the intersection of machine learning (ML) and neuromorphic computing has lead to advances in the development of physical neural networks (PNNs) [6]. These neural networks seek to harness the capabilities of various analog systems to improve both the speed and resource efficiency with which these models can be trained and deployed for inference.
@@
A promising area being explored for neuromorphic computing, and thus for the implementation of PNNs, is the field of photonics [5, 6]. Neuromorphic computing implementations based on photonics have several advantages such as higher energy efficiency, throughput, speed and parallelization capability [5]. Moreover, quantum photonics, particularly with continuous variable (CV) cluster states, offers outstanding scalability as much work has been done on scaling these systems to large sizes (thousands to millions entangled modes) for measurement-based quantum computing [7, 8, 9, 10]. One challenge has been the implementation of the nonlinearities, i.e, neuron activation, required for learning. Initial approaches have relied on physical nonlinearities such as opto-electronic conversion [11, 12, 13, 14, 15, 16, 17], the intrinsic nonlinearity within coupled semiconductor nanolaser arrays [18], or have utilized hybrid architectures based on optical-digital conversion [19, 20, 21]. On the quantum side, initial proposals utilized optical Kerr nonlinearities [22, 23] or photon-number-resolving measurements [24].
@@
But learning nonlinearity doesn’t require a physical one. Indeed, neuromorphic computing has recently been proposed [25] and demonstrated [26, 27] using classical linear wave scattering. These breakthroughs stemmed from the realization that optical nonlinearities in the field variables are not needed to implement learning [28] if one doesn’t exclusively encode the data in the fields but also uses the experimental parameters for that purpose.
@@
In this paper, we propose a different linear optics encoding: a multimode laser interferometer with added amplitude and phase displacements, see Fig.1. Indeed, a simple optical phase shift $\theta$ imparts field transformations, $E\cos\omega t\mapsto E\cos(\omega t+\theta)$, which are nonlinear in $\theta$, if linear in the field amplitude $E$. This approach is considerably simpler than the previous ones in terms of experimental implementation, does not utilize digital linear layers for postprocessing, and is also compatible with the constantly evolving state of the art of integrated photonics [29].
@@
It is interesting—in particular to set the stage for future extensions to the quantum regime—to deliberately adopt a quantum optical description of the optical circuit, even though it sits firmly within the classical border in this paper. Indeed, three classes of optical circuits can be defined: (i), classical linear optics (laser interferometers with photodetectors of dark current well above the single-photon level), (ii), second-order nonlinear optics yielding only two-wave mixing (quantum squeezers, e.g. undepleted-pump optical parametric oscillators) and, (iii), true three-wave mixing as well as third- and higher order nonlinear optics. While (ii) and (iii) both enter the realm of quantum optics, only (iii) involves nonlinear quantum evolution, a.k.a. non-Gaussian Wigner function gates and/or states of the quantum fields. This is why such operations—be they implemented via third-order nonlinearities or Fock state projection—were included in previous works on quantum neural networks [22, 24, 23]. Again, the gist of linear optics implementations [28, 25, 26, 27] is to eschew relying solely on field encoding, and hence the need for optical nonlinearities, in favor of optical parameter encoding.
@@
2 Laser interferometer neural network
2.1 Gaussian quantum formalism
As mentioned above, we adopt here the more general quantum optical formalism of qumodes in continuous-variable quantum information [30, 31, 32], for both the sake of generality and in anticipation of future extensions beyond the classical context of this work. Consider an $M$-qumode system [33] given by the vector of quadrature operators $Q_{j}=(a_{j}+a_{j}^{\dagger})/\sqrt{2}$ and $P_{j}=i(a_{j}^{\dagger}-a_{j})/\sqrt{2}$ of each qumode $j$, $a_{j}$ being the photon annihilation operator,
@@ 2.1 Gaussian quantum formalism
and the covariance matrix $\Sigma\in\mathbb{R}^{2M\times 2M}$ is given by
@@
$\Sigma$ $\mapsto\mathbf{S}\Sigma\mathbf{S}^{T}.$ (5)
The symplectic matrix representations of the single-mode phase shift $\mathbf{R}(\phi)$ and two-mode beamsplitter $\mathbf{B}(\theta)$ unitaries are respectively given by [31]
@@
2.2 Nonlinear learning with linear optics
@@
To approximate a fully connected linear layer present in a classical neural network we use an $M\times M$ interferometer containing beamsplitters of tunable transmissivity angle $\theta$ and fixed zero phase. To maintain consistency with classical neural networks’ fully connected layers, we employ beamsplitter operations between all M(M1)/2M(M-1)/2 pairs of modes, contrasting with conventional nearest-neighbor Mach-Zehnder interferometer (MZI) approaches [35, 36]. Specifically, given the mode indices $\{1,2,...,M\}$ the interferometer applies a beamsplitter operation between each pair of modes in the set of all two-element subsets denoted as
@@ 3.1 Gradient descent
3.2 Feasibility of in situ training
@@
To perform training in situ one will need a way of computing the gradients of the loss function with respect to network parameters using the physical system itself [6]. To this end many methods have been proposed for extracting gradients of PNN parameters such as equilibrium propagation for energy based models [39, 40, 41], scatter backpropagation for driven-dissipative nonlinear coupled-mode systems [42] and Hamiltonian echo backpropagation for lossless time-reversible Hamiltonian systems [43]. Our approach is compatible with well known single parameter shift methods which give exact gradients [44], physical backpropagation methods which have $O(1)$ complexity with respect to the number of parameters in the model [45] as well as stochastic simultaneous parameter shift methods which combine the experimental simplicity of the former with the efficiency of the latter [17].
@@
4 Numerical Experiments
Here we present results from numerical simulations of training our linear optical circuit to solve supervised regression and classification tasks using stochastic gradient descent. We utilized the Adam optimizer [46] implemented in PyTorch which uses automatic differentiation to compute the gradients [34]. The pseudocode for the training process is provided in the Supplemental Material [47]. Further details on the pseudocode can be found in [46, 48]. For all experiments we used a learning rate scheduler that decays the learning rate exponentially after each epoch. The parameters for all models were initialized randomly from the uniform distribution over the interval [1,1][-1,1]. All feature scaling is performed using the MinMaxScaler class from the scikit-learn Python library [49] with the exception of the handwritten digits dataset which was normalized by dividing the pixel intensities by 16. We did not find that regularization was necessary for our experiments but note that regularization of the displacement parameters, such as an $L_{1}$ or $L_{2}$ penalty, could be used to further improve energy efficiency as this is the only non-passive optical operation in the circuit.
4.1 Nonlinear regression
The first supervised learning problem we investigated was that of nonlinear regression. In regression problems, e.g. curve fitting, one attempts to learn a function which maps inputs to continuous outputs [50]. Our circuit architecture used a three mode circuit with three layers. The input scalar $x$ was repeatedly encoded in the coherent state amplitude of the three modes and in the phase angle of the phase shift operations on each mode in each subsequent layer. We fit four nonlinear functions: $\tanh(2x)$, $\sin(\pi x)$, $x^{2}$ and $x^{3}$. We took the expectation value of the $Q$ quadrature of the first mode as our output. When generating the training data we added a Gaussian noise term to the ground truth labels with mean zero and standard deviation $\sigma=0.05$. We generated 100 samples and used 50 for training and 50 for testing. The hyperparameters used for the regression training are provided in the Supplemental Material [47]. For our loss function we used the mean squared error function as used in previous works [22, 24].

### v4

@@ Abstract
We present a method for implementing an optical neural network using only linear optical resources, namely field displacement and interferometry applied to coherent states of light. The nonlinearity required for learning in a neural network is realized via an encoding of the input into phase shifts allowing for far more straightforward experimental implementation compared to previous proposals for, and demonstrations of, in situ inference. Beyond in situ inference, the method enables in situ training by utilizing established techniques like parameter shift methods or physical backpropagation to extract gradients directly from measurements of the linear optical circuit. We also investigate the effect of photon losses and find the model to be very resilient to these.
preprint: APS/123-QED
I Introduction
@@
Modern deep learning models have reached unprecedented sizes with recent large language models having billions of trainable parameters and taking months of time for pretraining on clusters of multiple graphics processing units (GPUs) [9, 43, 53]. The training and inference phases for these models require increasingly large amounts of energy due to the energy-efficiency limitations imposed by the modern silicon-based digital computers on which they are implemented, a phenomenon referred to as the ‘von Neumann bottleneck’. This bottleneck is a consequence of the separation of computing and memory into distinct units in the current von-Neuman computer architecture which necessitates costly data transfer between these units during computation [35, 12]. This challenge has motivated the development of neuromorphic computing, a field that seeks to construct hardware architectures inspired by the remarkable energy efficiency of the human brain [35, 12]. Consequently, the intersection of machine learning (ML) and neuromorphic computing has lead to advances in the development of physical neural networks (PNNs) [37]. These neural networks seek to harness the capabilities of various analog systems to improve both the speed and resource efficiency with which these models can be trained and deployed for inference.
@@
A promising area being explored for neuromorphic computing, and thus for the implementation of PNNs, is the field of photonics [12, 37]. Neuromorphic computing implementations based on photonics have several advantages such as higher energy efficiency, throughput, speed and parallelization capability [12]. However, a limitation of current approaches based on photonics is scalability and difficulty of implementing low power optical nonlinearities [12]. On the scalability front quantum photonics, particularly with continuous variable (CV) cluster states, can provide a remedy as much work has been done on scaling these systems to large sizes for measurement-based quantum computing schemes [Chen2014, Yoshikawa2016, Larsen2019, Asavanant2019]. The original proposal for a CV quantum neural network [27], as well as derivative work [5], utilized optical Kerr nonlinearities which are difficult to implement experimentally. An alternative proposal utilized repeat-until-success photon number resolving measurements on ancillary modes to introduce nonlinearity into the network [6]. However, the implementation of quantum optical nonlinearities is still a challenging task as in the classical optics case. One solution that has been explored in the classical optics approach is to utilize hybrid architectures based on optical-digital conversion [50, 55, 40] or opto-electronic conversion [49, 58, 18, 4, 11, 63, Bandyopadhyay2024]. Another approach uses the intrinsic nonlinearity within coupled semiconductor nanolaser arrays [25].
@@
However, learning nonlinearity should not be equated with physics nonlinearity. Indeed, recent breakthroughs in optical implementations [PT2024, 56, 64, 61] stemmed from the realization that optical nonlinearities in the field variables are not needed to implement learning [PT2024] if one doesn’t encode the data in them exclusively and, rather, uses the experimental parameters for that purpose: a simple optical phase shift $\theta$ imparts field transformations $\cos\theta,\sin\theta$ that are nonlinear in $\theta$ (though linear in the fields).
This has recently been proposed [56] and demonstrated [64, 61] using classical linear wave scattering.
@@
In this paper, we propose a different linear optics encoding: a multimode laser interferometer with added amplitude and phase displacements, see Fig.1. This approach is considerably simpler than the previous ones in terms of experimental implementation, does not utilize digital linear layers for postprocessing, and is also compatible with the current state of integrated photonics [Bogaerts2020].
@@
It is interesting—in particular to set the stage for future extensions to the quantum regime—to deliberately adopt a quantum optical description of the optical circuit, even though it sits firmly within the classical border in this paper. Indeed, three classes of optical circuits can be defined: (i), classical linear optics (laser interferometers with photodetectors of dark current well above the single-photon level), (ii), second-order nonlinear optics yielding only two-wave mixing (quantum squeezers, e.g. undepleted-pump optical parametric oscillators) and, (iii), true three-wave mixing as well as third- and higher order nonlinear optics. While (ii) and (iii) both enter the realm of quantum optics, only (iii) involves nonlinear quantum evolution, a.k.a. non-Gaussian Wigner function gates and/or states of the quantum fields. This is why such operations—be they implemented via third-order nonlinearities or Fock state projection—were included in previous works on quantum neural networks [27, 6, 5]. Again, the gist of linear optics implementations [PT2024, 56, 64, 61] is to eschew relying solely on field encoding, and hence the need for optical nonlinearities, in favor of optical parameter encoding.
@@
II Laser interferometer neural network
II.1 Gaussian quantum formalism
As mentioned above, we adopt here the more general quantum optical formalism of qumodes in continuous-variable quantum information [Braunstein2005a, Weedbrook2012, Pfister2019], for both the sake of generality and in anticipation of future extensions beyond the classical context of this work. Consider an $M$-qumode system [1] given by the vector of quadrature operators $Q_{j}=(a_{j}+a_{j}^{\dagger})/\sqrt{2}$ and $P_{j}=i(a_{j}^{\dagger}-a_{j})/\sqrt{2}$ of each qumode $j$, $a_{j}$ being the photon annihilation operator,
@@ II.1 Gaussian quantum formalism
and the covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{2M\times 2M}$ is given by
@@
$\mathbf{\Sigma}$ $\mapsto\mathbf{S}\mathbf{\Sigma}\mathbf{S}^{T}.$ (5)
The symplectic matrix representations of the single-mode phase shift $\mathbf{R}(\phi)$ and two-mode beamsplitter $\mathbf{B}(\theta)$ unitaries are respectively given by [Weedbrook2012]
@@
Algorithm 1 ONN training loop with Adam algorithm
1:Initial parameters $\theta_{0}$, initial learning rate $\gamma_{0}$, maximum number of epochs $T$, exponential decay rates $\beta_{1},\beta_{2}\in[0,1)$, batch size $B$, dataset $D=\{(x^{(1)},y_{\textrm{tar}}^{(1)}),\dots,(x^{(k)},y_{\textrm{tar}}^{(k)})\}$ of $k$ labeled training examples, learning rate decay rate $\chi$, small constant $\epsilon$ for numerical stability (typically set to $10^{-8}$)
[…]
5:while $t<T$ do
6:  $t\leftarrow t+1$
7:  while all data points in $D$ not yet sampled do $\triangleright$ An epoch is one pass over all data points
8:   Sample random batch of $B$ data points, $D_{B}=\{(x^{(1)},y_{\textrm{tar}}^{(1)}),\dots,(x^{(B)},y_{\textrm{tar}}^{(B)})\}$, without replacement from $D$
9:   Compute ONN predictions, $y=\{f(x^{(1)};\theta_{t-1}),\dots,f(x^{(B)};\theta_{t-1})\}$, from optical measurements
10:   $g_{t}\leftarrow\frac{1}{B}\sum_{i}\nabla_{\theta_{t-1}}\mathcal{L}(y^{(i)},y_{\textrm{tar}}^{(i)})$ $\triangleright$ Compute average Euclidean gradient over batch
11:   $m_{t}\leftarrow\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$ $\triangleright$ Update biased first moment estimate
12:   $v_{t}\leftarrow\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}\odot g_{t}$ $\triangleright$ Update biased second moment estimate. $\odot$ denotes Hadamard product.
13:   $\hat{m}_{t}\leftarrow\dfrac{m_{t}}{1-\beta_{1}^{t}}$ $\triangleright$ Correct bias in first moment estimate
14:   $\hat{v}_{t}\leftarrow\dfrac{v_{t}}{1-\beta_{2}^{t}}$ $\triangleright$ Correct bias in second moment estimate
15:   $\theta_{t}\leftarrow\theta_{t-1}-\gamma\dfrac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ $\triangleright$ Compute updated optical circuit parameters
16:  end while
17:  $\gamma_{t}\leftarrow\gamma_{0}\chi^{t}$ $\triangleright$ Decay learning rate on exponential schedule
18:end while
19:return $\theta_{t}$
II.2 Nonlinear learning with linear optics
@@
To approximate a fully connected linear layer present in a classical neural network we use an $M\times M$ interferometer containing beamsplitters of tunable transmissivity angle $\theta$ and fixed zero phase. To maintain consistency with classical neural networks’ fully connected layers, we employ beamsplitter operations between all M(M1)/2M(M-1)/2 pairs of modes, contrasting with conventional nearest-neighbor Mach-Zehnder interferometer (MZI) approaches [44, Clements2016]. Specifically, given the mode indices $\{1,2,...,M\}$ the interferometer applies a beamsplitter operation between each pair of modes in the set of all two-element subsets denoted as
@@ III.1 Gradient descent
III.2 Feasibility of in situ training
@@
To perform training in situ one will need a way of computing the gradients of the loss function with respect to network parameters using the physical system itself [37]. To this end many methods have been proposed for extracting gradients of PNN parameters such as equilibrium propagation for energy based models [46, 47, 57], scatter backpropagation for driven-dissipative nonlinear coupled-mode systems [13] and Hamiltonian echo backpropagation for lossless time-reversible Hamiltonian systems [34]. Here we present two well known methods for computing gradients in situ which are applicable to our optical circuit.
III.2.1 Parameter-shift method
For quantum circuits since our output is an expectation value of an observable, e.g. in our case one of the field quadratures, these gradients are also expectation values. The purely Gaussian nature of our circuit makes in situ training much more feasible than previous proposals [27, 6, 5] which use arbitrary numbers of non-Gaussian operations. For any purely Gaussian quantum circuit $U(\vec{\theta})$, $\vec{\theta}=(\theta_{i})_{i}^{T}$, it has been proven that the gradient elements $\partial_{\theta_{i}}\langle U^{\dagger}(\vec{\theta})OU(\vec{\theta})\rangle$ of the expectation values of the output observable $O$ with respect to circuit parameters $\theta_{i}$ can be computed via a simple shift rule of parameter $\theta_{i}$ [48]. Here, $O$ must be a low-degree polynomial in the field quadratures. Remarkably, these gradient elements can be obtained by using the same circuit that is used for inference, the only modification being that the value of $\theta_{i}$ is shifted by any nonzero constant twice. This result significantly simplifies the process of computing gradients for in situ training of Gaussian quantum circuits.
As a concrete example for the beamsplitter gate consider a circuit which applies a Gaussian unitary $G_{1}$ followed by a zero phase beamsplitter $BS(\theta)$ and another Gaussian unitary $G_{2}$ to a two-mode vacuum state input. The expectation value of the $Q$ quadrature is given by
[…]
Similarly in the case where the beamsplitter operation is replaced with a displacement operation $D(\alpha)$, where $\alpha=re^{i\phi}$, the parameter shift rules are given by
[…]
where $s$ is any nonzero real number. Similar rules are known for the squeezing and phase shift operations. These parameter shift rules, which are currently only known for Gaussian gates [7]—which is of course the case here—provide exact expressions for the gradients and thus provide an unbiased estimator of the partial derivative unlike methods based on finite differences which are only exact in the case of infinitesimal shifts [48, 52]. It has also been shown empirically in qubit based systems that the variance in the gradient estimate from using a finite number of shots still allows for good model performance even in the extreme case of only a few shots and in fact these estimators can be used to implement a form of gradient descent for which convergence guarantees have been proven [52]. The number of shots used to estimate the gradients can also be tuned as a hyperparameter and algorithms which select the number of shots in an adaptive and frugal manner while still achieving good performance have been developed [29]. The use of only Gaussian resources also makes a sample complexity analysis for estimating quadrature means straightforward. When using the the sample mean $\bar{q}$ to estimate the expectation value of the $Q$ quadrature the variance of the estimator is given by [15]
$\textrm{Var}(\bar{q})=\frac{\sigma^{2}}{n},$ (20)
where $n$ is the number of quadrature samples or shots and $\sigma^{2}$ is the variance of the quadrature which we know is equal to 1/2 since no squeezing is applied in the circuit and all inputs are coherent states. As the sample mean is an unbiased estimator of the population mean the variance is also equal to the mean squared error (MSE) and thus if one wants to estimate the $Q$ quadrature mean to within an MSE of $\epsilon$ one needs
$n=\frac{\sigma^{2}}{\epsilon}$ (21)
quadrature samples.
III.2.2 Simultaneous perturbation stochastic approximation
While the parameter shift rule given above provides a simple and unbiased estimator of the necessary gradients it’s computational complexity scales linearly with the number of parameters in the model. There do, however, exist methods for approximating the gradient that are based on similar principles of perturbing the function parameters but that perturb all parameters simultaneously. One example of such a method is simultaneous perturbation stochastic approximation (SPSA) [51]. This method was used to train the ONNs in [Bandyopadhyay2024, 60, 62].
The procedure consists of generating a random perturbation vector $\Delta\in\{+\epsilon,-\epsilon\}^{N}$, i.e. a vector of length $N$, where $N$ is the number of parameters in the model, whose components are independent and randomly distributed over the set $\{+\epsilon,-\epsilon\}$. Here $\epsilon$ is a hyperparameter that can be optimized and the distribution one uses to generate the values of $\Delta$ must satisfy certain conditions. The Bernoulli $\pm 1$ distribution with $p=0.5$, equivalent to the Rademacher distribution, is an example of such a distribution [51]. One may then compute
$\widehat{\frac{\partial f}{\partial\theta_{i}}}=\frac{f(x;\Theta+\Delta)-f(x;\Theta-\Delta)}{2c\Delta_{i}}$ (22)
where we use $\widehat{\frac{\partial f}{\partial\theta_{i}}}$ as this is a stochastic estimate of the gradient and $c$ is a small positive number which typically decreases as the optimization proceeds [51]. Under reasonably general conditions it can be proven this method and finite differences reach the same level of statistical accuracy given the same number of iterations even though SPSD is $N$ times more efficient and noisy [51].
@@ III.2.2 Simultaneous perturbation stochastic approximation
IV Numerical Experiments
Here we present results from numerical simulations of training our linear optical circuit to solve supervised regression and classification tasks using stochastic gradient descent. We utilized the Adam optimizer [28] implemented in PyTorch which uses automatic differentiation to compute the gradients [41]. The pseudocode for the training process is given in Algorithm 1. Further details on the pseudocode can be found in [28, 17]. For all experiments we used a learning rate scheduler that decays the learning rate exponentially after each epoch. The parameters for all models were initialized randomly from the uniform distribution over the interval [1,1][-1,1]. All feature scaling is performed using the MinMaxScaler class from the scikit-learn Python library [42] with the exception of the handwritten digits dataset which was normalized by dividing the pixel intensities by 16. We did not find that regularization was necessary for our experiments but note that regularization of the displacement parameters, such as an $L_{1}$ or $L_{2}$ penalty, could be used to further improve energy efficiency, as we show in Appendix B, as this is the only non-passive optical operation in the circuit.
Hyperparameter Value
Learning rate ($\gamma$) $10^{-2}$
Learning rate decay ($\chi$) $0.999$
11st moment estimates decay rate ($\beta_{1}$) $0.9$
22nd moment estimates decay rate ($\beta_{2}$) $0.999$
Number of epochs $1000$
Batch size $32$
Table 1: Hyperparameters used when training the optical circuit for regression tasks.
(a) XOR classification problem.
(b) Circles classification problem.
(c) Moons classification problem.
Figure 4: Training data used for binary classification problems. We set the standard deviation of the Gaussian noise to $\sigma=0.05$.
IV.1 Nonlinear regression
The first supervised learning problem we investigated was that of nonlinear regression. In regression problems, e.g. curve fitting, one attempts to learn a function which maps inputs to continuous outputs [24]. Our circuit architecture used a three mode circuit with three layers. The input scalar $x$ was repeatedly encoded in the coherent state amplitude of the three modes and in the phase angle of the phase shift operations on each mode in each subsequent layer. We fit four nonlinear functions: $\tanh(2x)$, $\sin(\pi x)$, $x^{2}$ and $x^{3}$. We took the expectation value of the $Q$ quadrature of the first mode as our output. When generating the training data we added a Gaussian noise term to the ground truth labels with mean zero and standard deviation $\sigma=0.05$. We generated 100 samples and used 50 for training and 50 for testing. The hyperparameters used for the regression training are shown in Table 1. For our loss function we used the mean squared error function as used in previous works [27, 6].